OpenStudy (firechild17):

Bonnie has a container in the shape of a rectangular pyramid. The formula for the surface area of the enclosed space is S = lw + 0.5Ph. Solve for P. P = S – lw – 0.5h P = S + lw + 0.5h P = The quantity S minus l times w all divided by 0.5 times h P = S divided by the quantity l times w plus 0.5 times h

OpenStudy (firechild17):

@alivejeremy

OpenStudy (phi):

you start with S = lw + 0.5Ph we want "P" by itself on one side. I would first "move" the term lw to the other side lw (or l*w or l times w) can be thought of as some number if you subtract it from the right side (and the left side to keep things equal) you have S- lw = lw - lw + 0.5Ph any idea what lw - lw simplifies to ?

OpenStudy (michaelbp):

are those the answers

OpenStudy (firechild17):

thats confusing...

OpenStudy (firechild17):

thats confusing @phi

OpenStudy (phi):

OK, if it's confusing, that means you are missing an idea to figure out what idea you need to learn, do you know how to solve for P with S= 2+P ?

OpenStudy (firechild17):

i think

OpenStudy (phi):

how ?

OpenStudy (firechild17):

u cant do it

OpenStudy (will.h):

Yes we can

OpenStudy (phi):

solve means "have P on one side, by itself"

OpenStudy (will.h):

How do we isolate p

OpenStudy (firechild17):

how? im like so lost

OpenStudy (firechild17):

you have to subtract 2 from both sides

OpenStudy (phi):

ok how about 5= 2+ P can you get "P" by itself ?

OpenStudy (will.h):

S= 2 + p by phi Solve for p P is having 2 added to it So we can get rid of 2 by reversing the operation. What's the opposite of adding?

OpenStudy (firechild17):

yes @phi subtacting @will.h

OpenStudy (will.h):

Great So subtract 2 from both sides S - 2 = 2 - 2 + p The right side can be simplified 2-2 is zero so that would leave only p S-2 = p Your original problem can be solved just like this

OpenStudy (phi):

Here are the two ideas you need to solve for P in 5= 2+p 1) 2 - 2 = 0 (which makes it "go away") so if on the right side we do 2- 2 + p the 2-2 becomes 0 and 0+p= p because adding 0 does not change anything 2) if we do -2 on one side, we *must* do the same on the other side, to keep things equal

OpenStudy (phi):

does that part make sense ?

OpenStudy (firechild17):

ok @phi @Will.H

OpenStudy (phi):

as a test, can you "solve for p" in 6= 1+P ?

OpenStudy (firechild17):



OpenStudy (firechild17):

like that?

OpenStudy (phi):

ok. next idea: anything minus itself is zero that is an important idea.

OpenStudy (phi):

we use that idea in 5= x+P how do we "solve for P" ? what do we subtract from both sides ?

OpenStudy (phi):

is this part confusing ?

OpenStudy (firechild17):

we subtract x

OpenStudy (firechild17):

but i dont know how

OpenStudy (phi):

with letters it's easier than with numbers. you write -x on both sides which means "subtract x" example: 5= x+P 5-x = x-x+P

OpenStudy (firechild17):

ok

OpenStudy (phi):

we can't "simplify" 5-x (it stays that way) but on the right side, we know x-x is 0 because anything (which includes "x") minus itself is 0 so we can simplify it to 5-x = 0+P and (I think you know) adding 0 can be ignored (or rather 0+ whatever = whatever) so we get 5-x= P

OpenStudy (firechild17):

ok and how does this help with my math problem?

OpenStudy (phi):

what to try solving for P in S= lw + P ? (lw means l*w, but we can subtract lw from both sides)

OpenStudy (phi):

*want (not what)

OpenStudy (firechild17):

yes so p=lw-s

OpenStudy (phi):

yes. so we are getting there!

OpenStudy (firechild17):

ok so whats next?

OpenStudy (phi):

the original problem is S = lw + 0.5Ph we know how to subtract lw from both sides to get 0.5Ph = S-lw (I switched the left and right sides, but we can do that)

OpenStudy (firechild17):

ok so now im confused

OpenStudy (phi):

let me do it slower S = lw + 0.5Ph add -lw to both sides (we write -lw on both sides) S-lw = lw - lw + 0.5Ph simplify lw - lw to 0 S-lw = 0 + 0.5Ph next simplify 0+0.5Ph to 0.5Ph (adding 0 is ignored) S-lw= 0.5Ph ok ?

OpenStudy (firechild17):

ok

OpenStudy (phi):

and S-lw= 0.5Ph can also be written 0.5Ph = S-lw

OpenStudy (firechild17):

so the answer is A?

OpenStudy (phi):

I assume you know if A=B then B=A i.e. in an equation, we can write it either way. usually, we write it so the variable we are "solving for" is on the left side so before continuing, we write S-lw= 0.5Ph as 0.5Ph = S-lw

OpenStudy (phi):

the next idea this one is as important as anything minus itself is zero IDEA: anything divided by itself is one examples: 2/2 is 1 5/5 is 1 x/x is 1 xy/ xy = 1 and so on

OpenStudy (firechild17):

ok so its b?

OpenStudy (phi):

we have 0.5Ph on the left side that means 0.5 * P *h (they are being multiplied) we can change the order of multiply (example 2*3 is the same as 3*2) in this case, write it as 0.5*h * P now we divide this by 0.5*h, like this \[ \frac{0.5\cdot h \cdot P}{0.5\cdot h} \] or (if you know how fractions work) \[ \frac{0.5\cdot h }{0.5\cdot h} P\]

OpenStudy (phi):

notice we have 0.5h divided by itself that means we get 1 and the left side 0.5hP/(0.5h) = 1*P= P (1 times anything is the anything) but to keep things equal we have to divide the other side by 0.5h

OpenStudy (phi):

in other words we do \[ \frac{0.5Ph}{0.5h}= \frac{S-lw}{0.5h} \\ P= \frac{S-lw}{0.5h} \] notice I did not rearrange the top from 0.5Ph to 0.5h P because I know the 0.5 divides with the 0.5 in the bottom and ditto for the h

OpenStudy (firechild17):

ok

OpenStudy (firechild17):

so whats the answer? now im confused

OpenStudy (phi):

read the last post carefully

OpenStudy (firechild17):

Its C?

OpenStudy (phi):

which of your choices matches what I posted ? yes, choice C (if you wrote it as an equation instead of in words)

OpenStudy (firechild17):

ok